notmeta:

15x + 17 1/3 = (23.5-45) 3.45 idk how to do this :(

notmeta:

idk how to even start tbh (ik this is easy for some people but i honestly dont understand it)

Manny300303199:

are we supposed to graph it

notmeta:

no just solve :/

notmeta:

"solve for x"

notmeta:

15x + 17 1/3 = (23.5-45) 3.45 i legit think its not solvable

Manny300303199:

I dont think it is

notmeta:

hm :/

Manny300303199:

there is no other x

notmeta:

shi i forgot the x at the end 15x + 17 1/3 = (23.5-45) 3.45x

EN0KI:

\[15x + 17\frac{ 1 }{ 3 } = (23.5-45) 3.45x \] Err I may be wrong but this is ur equation right?

EN0KI:

Like that's how its set up?

notmeta:

yes

Manny300303199:

@en0ki wrote:
\[15x + 17\frac{ 1 }{ 3 } = (23.5-45) 3.45x \] Err I may be wrong but this is ur equation right?
thast what i was think x at the end of 3.45

Manny300303199:

thinking*

notmeta:

thats the problem

EN0KI:

Mkay well you're going to want to turn the mixed number (17 1/3) into a fraction. So 17 times 3= 51+1= 52/3 So then it becomes:\[15x+\frac{ 52 }{ 3 }=\left(23.5-45\right)3.45x\]

notmeta:

is it the same as pemdas with solving the things in the brackets first?

EN0KI:

And \[23.5-45 \times 3.45x = -74.175x\] Then your equation becomes: \[15x+\frac{ 52 }{ 3 }=-74.175x\] Move 52/3 to the right, \[15x=-74.175x-17.3333 \] and so on 74.175x to the left \[89.175=-17.3333\] Divide both sides by 89.175x And then you get x=-0.19437 I believe

EN0KI:

@notmeta wrote:
is it the same as pemdas with solving the things in the brackets first?
Slightly. Except you're not starting with parenthesis, just fraction first then parenthesis and so forth.

notmeta:

thank you fo you're help.